OpenStudy (anonymous):

What are the odds in favor of drawing a card higher than a jack from an ordinary deck of cards? _____to____

OpenStudy (dumbcow):

there are 4 Q, 4K, and 4A 12 cards higher than J in deck of 52 P(drawing higher) = 12/52 = 3/13

OpenStudy (dumbcow):

odds = P/1-P = (3/13)/(10/13) = (3/13)*(13/10) = 3/10 or 3:10